# Lookahead Q-Cache: Achieving More Consistent KV Cache Eviction via Pseudo Query

### v1

@@ Abstract
Yixuan Wang [^2] Equal contribution. Shiyu Ji [^2] Equal contribution. Yijun Liu Yuzhuang Xu Yang XuQingfu ZhuWanxiang Che [^1] Corresponding author. Research Center for Social Computing and Interactive Robotics, Harbin Institute of Technology, China {yixuanwang,car}@ir.hit.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong capabilities in long-sequence text modeling tasks (Liu et al., 2025a, b) such as code generation (Guo et al., 2024a; Hui et al., 2024), document summarization (Liu et al., 2024c), and mathematical reasoning (Chen et al., 2025). To improve inference efficiency during the decoding stage, LLMs leverage key-value cache (KV-Cache) to reduce redundant computations and accelerate inference. However, increasing text lengths lead to a substantial rise in KV cache memory usage, introducing considerable obstacles (Shi et al., 2024; Li et al., 2024a; Yuan et al., 2024) to efficient model deployment. Recent work (Liu et al., 2024d; Sun et al., 2024) has focused on lightweighting the KV cache in long-context scenarios.
@@
As a straightforward and effective compression method, KV cache eviction has attracted widespread attention (Ge et al., 2024; Liu et al., 2023; Tang et al., 2024) from researchers. The method improves decoding efficiency by performing token-level pruning on the prefilling-stage KV cache. Following the observations of Liu et al. (2023); Zhang et al. (2023), the majority of studies adopt cumulative attention scores as the criterion for token pruning. Compared to direct accumulation, SnapKV (Li et al., 2024b) achieves more accurate KV cache importance estimation by leveraging an observation window. In addition, some studies (Cai et al., 2024; Feng et al., 2024) aim to allocate finer-grained budgets across layers or attention heads to achieve higher compression rates.
@@ 2.1 KV Cache Eviction
KV Cache eviction aims to prune redundant KV cache entries from the prefill stage at the token level, improving decoding speed and alleviating memory usage without compromising decoding performance. Previous work primarily builds on Liu et al.; Zhang et al.’s (2023; 2023) findings, utilizing the attention mechanism as an inherent evaluation criterion to assess the importance of KVCache entries and guide their eviction. However, using global attention scores often lacks specificity. Li et al. (2024b) discover that the suffix of the input window exhibits behavior more consistent with the generation phase. By maintaining a local window to guide KVCache eviction, lower performance degradation is achieved.
Due to conflicts between obtaining attention scores and existing acceleration techniques (such as FlashAttention (Dao et al., 2022; Dao, 2024)), some methods focus on identifying alternative importance evaluation metrics. Devoto et al. (2024) identify a correlation between the norm of the key values and their importance, and propose an L2 norm-based KV cache eviction method that is compatible with existing acceleration frameworks. Building on this, Guo et al. (2024b) also explore the norm of the value vectors and achieve more precise KV Cache eviction by combining the attention mechanism with the L2 norm of Value caches. Despite alleviating performance degradation, existing KV Cache eviction strategies continue to face the inconsistency problem in compressed inference illustrated in Figure 1, thereby constraining the upper bound of eviction performance.
@@ 2.2 KV Cache Compression
In addition to direct eviction, some studies (Wan et al., 2024) focus on merging similar KV cache entries to improve efficiency. To preserve as much information from the prefill stage as possible, some methods (Bolya et al., 2023; Zhang et al., 2024) build upon KV cache eviction and further explore token-level cache fusion, leading to improved task performance after eviction. Going further, with appropriate processing and training (Sun et al., 2024; Liu et al., 2024b; Zuhri et al., 2024; Ma et al., 2024), KV cache entries across layers can be merged and compressed to reduce memory consumption, offering additional efficiency gains.
Moreover, dimensionality reduction of the KV cache Shazeer (2019) serves as an effective approach to improve efficiency. The most straightforward compression method is to reduce the precision bit-width (He et al., 2024) of the KV Cache. Liu et al. (2024d) propose separate quantization methods based on channels and tokens to address the distinct distribution characteristics of the key and value, achieving nearly lossless 2-bit quantization. Additionally, some methods employ low-rank techniques for compression. GQA (Ainslie et al., 2023) achieves significant compression by sharing keys and values across multiple groups of queries, and has been widely adopted in various applications (Grattafiori et al., 2024). Multi-Head Latent Attention (MLA) (Liu et al., 2024a) of DeepSeek significantly reduces KV cache size through low-rank compression and decoupled RoPE, achieving efficient inference while maintaining model performance.
@@ 3 Observasions
As shown in Figure 2, we conduct evaluations on the GovReport dataset (Huang et al., 2021) for the summarization task. By analyzing the figure, we obtain several insightful findings.
@@
Similar to the conclusion of Li et al. (2024b), we find that local windows near the end of the input achieve higher recall rates compared to other input windows. However, once tokens from the output portion appear in the observation window (as indicated by the red vertical line in the figure), the recall rate of the window experiences a sudden increase. Although the recall rate gradually increases as the window shifts forward, just 1-2 tokens from the generation phase are sufficient to bring a substantial improvement. This phenomenon indicates a significant discrepancy between input and output queries and suggests that existing eviction methods still have substantial room for improvement.
@@ 4.2 Lookahead Query Construction
After obtaining the KV cache during the pre-filling stage, we first employ certain KV cache selection strategies to construct lookahead queries. To balance efficiency, we adopt KV cache eviction methods (such as StreamingLLM (Xiao et al., 2024) and snapKV (Li et al., 2024b)) with a low budget to generate a specified number of response tokens at a lower cost. Considering that the primary bottleneck for model generation in long-text scenarios lies in the pre-filling stage, a few additional decoding steps introduce minimal impact on the overall latency. We further analyze the latency introduced by the Lookahead Q-Cache in Section 6.3.
@@ 4.3 KV-Cache Re-eviction
where $W$ represents the local context window, similar to that employed in SnapKV (Li et al., 2024b). We denote this integrated method as LAQ++.
@@
Mistral-7B-v0.2-Instruct Full KV 26.90 33.14 49.26 42.77 27.35 18.77 32.95 24.21 27.13 71.00 86.23 42.72 2.75 86.98 57.12 54.51 42.74
@@
Llama3.1-8B-Instruct Full KV 31.85 15.55 28.17 29.93 22.98 18.20 34.39 23.79 27.12 72.50 92.14 43.66 8.37 97.59 65.05 54.78 41.63
@@ 5.1 Setup
Following prior works (Li et al., 2024b; Cai et al., 2024), we evaluate the proposed method using the LongBench benchmark (Bai et al., 2023) and the Needle-in-a-Haystack test (Kamradt, 2023). For LongBench, we adopt three KV cache budget settings: 128, 256, and 512. To comprehensively evaluate the generalization ability of the proposed method, we conduct experiments on Mistral-7B-v0.2-Instruct (Jiang et al., 2023), Llama3.1-8B-Instruct (Grattafiori et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2025). Due to space limitations, additional experimental results are presented in Appendix A.
@@
To highlight the effectiveness of our proposed method, we select three commonly used and strong KV-Cache eviction strategies as baselines: (1) H2O (Zhang et al., 2023) evaluates and evicts existing KV cache entries based on cumulative attention scores. (2) SnapKV (Li et al., 2024b) utilizes a local observation window to achieve more accurate importance estimation and employs average pooling to retain a more coherent KV Cache. (3) PyramidKV (Cai et al., 2024) leverages the characteristics of attention distribution across layers by allocating different budgets to the KV cache at shallow and deep layers, thereby achieving lower performance degradation under low-budget settings.
@@
As shown in Table 1, we evaluate the performance of various methods across multiple models on the LongBench benchmark. Overall, the proposed methods achieve a significant improvement of 1$\sim$4 percentage points over existing KV cache eviction strategies across different models and budget settings. By leveraging a pre-fetched lookahead Q-Cache, LAQ is able to make cache selection decisions that are more aligned with those in the actual inference stage. Furthermore, by integrating local observation windows, LAQ++ achieves an orthogonal improvement.
@@
As shown in Table 2, we conduct experiments on various models and budgets under a 32K context length. With a complete KV cache available, the model demonstrates strong capability in successfully completing the retrieval task. However, after evicting a certain portion of the KV cache, all methods experience some degree of performance degradation. The retrieval experiment results exhibit less variability and are more intuitive compared to task performance metrics. Although PyramidKV achieves significant advantages among baseline methods through dynamic budgeting, LAQ++ attains near-lossless performance by leveraging a more consistent Q-Cache mechanism. As discussed above, our proposed method can also be further enhanced by integrating dynamic budgeting. A more intuitive example is shown in Figure 4, where compared to SnapKV, LAQ++ can recall nearly all “needles” from a 32K context using only a 128 budget. Further experimental results can be found in Appendix B.
@@ 7 Conclusion
Appendix A Full Evaluation on LongBench
@@
Mistral-7B-v0.2-Instruct Full KV 26.90 33.14 49.26 42.77 27.35 18.77 32.95 24.21 27.13 71.00 86.23 42.72 2.75 86.98 57.12 54.51 42.74
@@
Llama3.1-8B-Instruct Full KV 31.85 15.55 28.17 29.93 22.98 18.20 34.39 23.79 27.12 72.50 92.14 43.66 8.37 97.59 65.05 54.78 41.63
@@
Qwen2.5-7B-Instruct Full KV 17.07 43.76 52.61 57.70 47.13 29.85 32.00 23.61 23.95 39.50 87.65 40.08 8.50 100.00 6.62 9.79 38.74
@@
Llama3-8B-Instruct Full KV 25.56 32.27 39.71 43.56 35.09 21.18 28.71 23.26 26.64 73.50 90.48 42.33 4.80 69.25 59.29 54.05 41.86
@@
Appendix B Evaluation on Needle-in-a-Haystack

### v2

@@ Abstract
Yixuan Wang [^2] Equal contribution.  Shiyu Ji [^2] Equal contribution.  Yijun Liu  Yuzhuang Xu Yang XuQingfu ZhuWanxiang Che [^1] Corresponding author. Research Center for Social Computing and Interactive Robotics, Harbin Institute of Technology, China {yixuanwang,car}@ir.hit.edu.cn  https://github.com/noforit/Lookahead_Q-Cache
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong capabilities in long-sequence text modeling tasks (liu2025comprehensive; liu2025survey) such as code generation (guo2024deepseek; hui2024qwen2), document summarization (liu2024sumsurvey), and mathematical reasoning (chen2025towards). To improve inference efficiency during the decoding stage, LLMs leverage key-value cache (KV-Cache) to reduce redundant computations and accelerate inference. However, increasing text lengths lead to a substantial rise in KV cache memory usage, introducing considerable obstacles (shi2024keep; li2024survey; yuan2024kv) to efficient model deployment. Recent work (liu2024kivi; sun2024you) has focused on lightweighting the KV cache in long-context scenarios.
@@
As a straightforward and effective compression method, KV cache eviction has attracted widespread attention (ge2023model; liu2023scissorhands; tang2024quest) from researchers. The method improves decoding efficiency by performing token-level pruning on the prefilling-stage KV cache. Following the observations of liu2023scissorhands; zhang2023h2o, the majority of studies adopt cumulative attention scores as the criterion for token pruning. Compared to direct accumulation, SnapKV (li2024snapkv) achieves more accurate KV cache importance estimation by leveraging an observation window. In addition, some studies (cai2024pyramidkv; feng2024ada) aim to allocate finer-grained budgets across layers or attention heads to achieve higher compression rates.
@@ 2.1 KV Cache Eviction
KV Cache eviction aims to prune redundant KV cache entries from the prefill stage at the token level, improving decoding speed and alleviating memory usage without compromising decoding performance. Previous work primarily builds on liu2023scissorhands; zhang2023h2o’s (liu2023scissorhands; zhang2023h2o) findings, utilizing the attention mechanism as an inherent evaluation criterion to assess the importance of KVCache entries and guide their eviction. However, using global attention scores often lacks specificity. li2024snapkv discover that the suffix of the input window exhibits behavior more consistent with the generation phase. By maintaining a local window to guide KVCache eviction, lower performance degradation is achieved.
Due to conflicts between obtaining attention scores and existing acceleration techniques (such as FlashAttention (dao2022flashattention; dao2023flashattention)), some methods focus on identifying alternative importance evaluation metrics. devoto2024simple identify a correlation between the norm of the key values and their importance, and propose an L2 norm-based KV cache eviction method that is compatible with existing acceleration frameworks. Building on this, guo2024attention also explore the norm of the value vectors and achieve more precise KV Cache eviction by combining the attention mechanism with the L2 norm of Value caches. Despite alleviating performance degradation, existing KV Cache eviction strategies continue to face the inconsistency problem in compressed inference illustrated in Figure 1, thereby constraining the upper bound of eviction performance.
@@ 2.2 KV Cache Compression
In addition to direct eviction, some studies (wan2024d2o) focus on merging similar KV cache entries to improve efficiency. To preserve as much information from the prefill stage as possible, some methods (bolya2022token; zhang2024cam) build upon KV cache eviction and further explore token-level cache fusion, leading to improved task performance after eviction. Going further, with appropriate processing and training (sun2024you; liu2024minicache; zuhri2024mlkv; ma2024compressing), KV cache entries across layers can be merged and compressed to reduce memory consumption, offering additional efficiency gains.
Moreover, dimensionality reduction of the KV cache shazeer2019fast serves as an effective approach to improve efficiency. The most straightforward compression method is to reduce the precision bit-width (he2024zipcache) of the KV Cache. liu2024kivi propose separate quantization methods based on channels and tokens to address the distinct distribution characteristics of the key and value, achieving nearly lossless 2-bit quantization. Additionally, some methods employ low-rank techniques for compression. GQA (ainslie2023gqa) achieves significant compression by sharing keys and values across multiple groups of queries, and has been widely adopted in various applications (grattafiori2024llama). Multi-Head Latent Attention (MLA) (liu2024deepseek) of DeepSeek significantly reduces KV cache size through low-rank compression and decoupled RoPE, achieving efficient inference while maintaining model performance.
@@ 3 Observasions
As shown in Figure 2, we conduct evaluations on the GovReport dataset (huang2021efficient) for the summarization task. By analyzing the figure, we obtain several insightful findings.
@@
Similar to the conclusion of li2024snapkv, we find that local windows near the end of the input achieve higher recall rates compared to other input windows. However, once tokens from the output portion appear in the observation window (as indicated by the red vertical line in the figure), the recall rate of the window experiences a sudden increase. Although the recall rate gradually increases as the window shifts forward, just 1-2 tokens from the generation phase are sufficient to bring a substantial improvement. This phenomenon indicates a significant discrepancy between input and output queries and suggests that existing eviction methods still have substantial room for improvement.
@@ 4.2 Lookahead Query Construction
After obtaining the KV cache during the pre-filling stage, we first employ certain KV cache selection strategies to construct lookahead queries. To balance efficiency, we adopt KV cache eviction methods (such as StreamingLLM (xiao2023efficient) and snapKV (li2024snapkv)) with a low budget to generate a specified number of response tokens at a lower cost. Considering that the primary bottleneck for model generation in long-text scenarios lies in the pre-filling stage, a few additional decoding steps introduce minimal impact on the overall latency. We further analyze the latency introduced by the Lookahead Q-Cache in Section 6.3.
@@ 4.3 KV-Cache Re-eviction
where $W$ represents the local context window, similar to that employed in SnapKV (li2024snapkv). We denote this integrated method as LAQ++.
@@
Mistral-7B-v0.2-Instruct
Full KV 26.90 33.14 49.26 42.77 27.35 18.77 32.95 24.21 27.13 71.00 86.23 42.72 2.75 86.98 57.12 54.51 42.74
@@
Llama3.1-8B-Instruct
Full KV 31.85 15.55 28.17 29.93 22.98 18.20 34.39 23.79 27.12 72.50 92.14 43.66 8.37 97.59 65.05 54.78 41.63
@@ 5.1 Setup
Following prior works (li2024snapkv; cai2024pyramidkv), we evaluate the proposed method using the LongBench benchmark (bai2023longbench) and the Needle-in-a-Haystack test (kamradt2023needle). For LongBench, we adopt three KV cache budget settings: 128, 256, and 512. To comprehensively evaluate the generalization ability of the proposed method, we conduct experiments on Mistral-7B-v0.2-Instruct (jiang2023mistral7b), Llama3.1-8B-Instruct (grattafiori2024llama), and Qwen2.5-7B-Instruct (yang2025qwen2). Due to space limitations, additional experimental results are presented in Appendix D.
@@
To highlight the effectiveness of our proposed method, we select three commonly used and strong KV-Cache eviction strategies as baselines: (1) H2O (zhang2023h2o) evaluates and evicts existing KV cache entries based on cumulative attention scores. (2) SnapKV (li2024snapkv) utilizes a local observation window to achieve more accurate importance estimation and employs average pooling to retain a more coherent KV Cache. (3) PyramidKV (cai2024pyramidkv) leverages the characteristics of attention distribution across layers by allocating different budgets to the KV cache at shallow and deep layers, thereby achieving lower performance degradation under low-budget settings.
@@
[…]
@@
As shown in Table 2, we conduct experiments on various models and budgets under a 32K context length. With a complete KV cache available, the model demonstrates strong capability in successfully completing the retrieval task. However, after evicting a certain portion of the KV cache, all methods experience some degree of performance degradation. The retrieval experiment results exhibit less variability and are more intuitive compared to task performance metrics. Although PyramidKV achieves significant advantages among baseline methods through dynamic budgeting, LAQ++ attains near-lossless performance by leveraging a more consistent Q-Cache mechanism. As discussed above, our proposed method can also be further enhanced by integrating dynamic budgeting. A more intuitive example is shown in Figure 4, where compared to SnapKV, LAQ++ can recall nearly all “needles” from a 32K context using only a 128 budget. Further experimental results can be found in Appendix E.
@@ 7 Conclusion
Acknowledge
We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC) via grant 62236004, 62206078, 62441603 and 62476073.
Appendix A Detailed Latency Comparison
To provide a more intuitive comparison of our method with other baselines, such as SnapKV, we conducted additional latency analysis experiments. These experiments were performed on a single A100-80GB GPU. We calculated the average wall-clock time (in seconds) for each method at different stages, with a fixed output length of 256 tokens and varying input lengths.
Table 3: Average wall-clock time (in seconds) for different methods. The output length is fixed at 256 tokens.
Method Pre-filling Stage 2 Decoding
INPUT 16K
Full KV 1.46615 - 12.38489
StreamingLLM 1.49165 - 8.79209
H2O 1.47285 - 8.62985
SnapKV 1.48718 - 8.64949
LAQ++ (2) 1.46968 0.12636 8.75256
LAQ++ (8) 1.46952 0.31302 8.61714
INPUT 32K
Full KV 3.58076 - 18.63942
StreamingLLM 3.65054 - 8.73857
H2O 3.58777 - 8.75228
SnapKV 3.61811 - 8.61096
LAQ++ (2) 3.59406 0.22643 8.83214
LAQ++ (8) 3.59088 0.41255 8.64124
As shown in the Table 3, the additional Stage 2 in our method incurs a minimal time overhead, typically less than 5%. This minor increase in latency is a small trade-off for the significant performance enhancements in the eviction algorithm, particularly under low-budget conditions (see Table 1 in the main text and Tables 6 and 7 in the appendix). This demonstrates the practical value of our approach.
Appendix B Analysis of Forward Templates
To investigate the impact of different forward templates, we conducted experiments using templates such as "The answer is" and "I have found the answer!". These templates were designed to provide a fixed semantic guide without requiring additional decode computation. The goal was to determine if the performance gains observed in our method, LAQ++, were due to semantic guidance rather than the query vector information derived from lookahead. The Table 4 shows the experimental results on the LongBench benchmark with a 512 KV-Cache budget for two different base models: Mistral-7B-v2.0-Instruct and Llama3.1-8B-Instruct.
Table 4: LongBench scores for different forward templates with a 512 KV-Cache budget.
Base Model Method Score
Mistral Full KV 42.74
Streaming 31.63
H2O 37.38
SnapKV 40.51
"The answer is" 38.20
"I have found the answer!" 37.79
LAQ++ 41.70
Llama Full KV 41.63
Streaming 34.61
H2O 35.72
SnapKV 38.87
"The answer is" 36.86
"I have found the answer!" 36.59
LAQ++ 40.39
Unfortunately, our findings indicate that these simple, fixed templates do not yield substantial performance improvements compared to other baselines. The scores for "The answer is" and "I have found the answer!" are significantly lower than those of our LAQ++ method and even SnapKV. This suggests that the performance gains from LAQ++ are likely not due to simple semantic guidance provided by the text, but rather from the pseudo vector information that is more effectively aligned with the output distribution. Consequently, constructing the Q-Cache through a lookahead mechanism remains a crucial and necessary component of our approach.
Appendix C RULER Task Evaluation
Beyond the standard LongBench and "needle in a haystack" benchmarks, we have included a more challenging evaluation using the RULER task to assess the long-context capabilities of our model. The experimental setup is consistent with the methodology described in Section 5.1. The following Table 5 presents the experimental results for the Llama 3.1-8B-Instruct model:
Table 5: RULER Task Performance on Llama 3.1-8B-Instruct
Method Budget Avg.
FullKV 32K 89.28
StreamingLLM 128 12.19
H2O 128 14.40
SnapKV 128 36.87
LAQ++ 128 51.57
StreamingLLM 256 13.31
H2O 256 25.25
SnapKV 256 51.58
LAQ++ 256 62.81
StreamingLLM 512 14.83
H2O 512 36.44
SnapKV 512 60.89
LAQ++ 512 69.83
As demonstrated by the results, the RULER task reveals a more significant performance differentiation among the various methods, particularly with a low KV-Cache budget. The proposed method, LAQ++, shows a nearly 10-point improvement over SnapKV when evaluated on this task.
Appendix D Full Evaluation on LongBench
@@
Mistral-7B-v0.2-Instruct
Full KV 26.90 33.14 49.26 42.77 27.35 18.77 32.95 24.21 27.13 71.00 86.23 42.72 2.75 86.98 57.12 54.51 42.74
@@
Llama3.1-8B-Instruct
Full KV 31.85 15.55 28.17 29.93 22.98 18.20 34.39 23.79 27.12 72.50 92.14 43.66 8.37 97.59 65.05 54.78 41.63
@@
Qwen2.5-7B-Instruct
Full KV 17.07 43.76 52.61 57.70 47.13 29.85 32.00 23.61 23.95 39.50 87.65 40.08 8.50 100.00 6.62 9.79 38.74
@@
Llama3-8B-Instruct
Full KV 25.56 32.27 39.71 43.56 35.09 21.18 28.71 23.26 26.64 73.50 90.48 42.33 4.80 69.25 59.29 54.05 41.86
@@
Appendix E Evaluation on Needle-in-a-Haystack